The diffraction bands are of variable thickness,
whereas interference bands are equidistant'.
Discuss the validity of this statement with
reference to the diffraction pattern of a straight
edge.
1
Expert's answer
2019-05-30T09:07:23-0400

The intensity profile can be calculated using the Fraunhofer diffraction equation as


"I(θ)=I_0\sin {c^2 \frac { d\pi}{\lambda}\sin θ}"

where I(θ) is the intensity at a given angle,

I0 is the original intensity


The distance between fringes is given by


"y=\frac { x \lambda }{d}"

where x is the distance to the screen,

d is the is the distance between the slits


In this way, the diffraction bands are of variable thickness because  all diffraction bands are of different intensities.